Find and correct the errors in the following mathematical statements.
(2x)^2 +5x = 4x+5x = 9x

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

LHS = (2x)^2+5x = 4x^2+5x≠ 9x = RHS
The correct statement is(2x)^2+5x = 4x^2+5x

Thanks